OpenStudy (anonymous):

What are the minimum, first quartile, median, third quartile and maximum of the data set? 60,50,130,200,180,150,100,140 answer choices are A: min-50, 1Q-80, Median-135, 3Q-165, Max-200 B: min-50, 1Q-107.5, Med-150, 3Q-182.5, Max-200 C: min-50, 1Q-80, Med-135, 3Q- 182.5, Max-200<---my answer choice D: Min-50, 1Q-65, Med-150, 3Q-165, Max-200 before this I solved it on my paper my answers were: Min-50 1Q-60 Med-135 3Q-180 Max-200

OpenStudy (texaschic101):

put the numbers in numerical order... 50,60,100,130,140,150,180,200 median (middle number) = (140 + 130)/2 = 270/2 = 135 first quartile -- (60 + 100)/2 = 160/2 = 80 third quartile -- (150 + 180)/2 = 330/2 = 165 minimum -- 50 maximum -- 200 I got the answer as A.

OpenStudy (anonymous):

Oh I see how you did it, I guess I did wrong, Thank you for explaining I don't look forward for answer I would want to learn how to do it. Thanks again ;)

OpenStudy (texaschic101):

your quite welcome.....they can sometimes be a little confusing :)

OpenStudy (anonymous):

oh yes they can.